OpenStudy (anonymous):

Solve by completing the square: x - 1/x = 1

OpenStudy (anonymous):

is it this ? (x-1)/x = 1 or x-(1/x) =1

OpenStudy (anonymous):

It's 

OpenStudy (anonymous):

u will have to apply quadratic equation

OpenStudy (anonymous):

so first we will find the common denominator on the left side to get (x^2 -1)/x=1 and then let us multiply by x on both sides to get x^2 -1 =x now let us bring the x over to the other side to get a quadratic equation which is x^2 -x -1 =0 now to complete the square we look at the middle number which is -1 divide by 2 then square it to get 1/4 so let us add 1/4 and subtract 1/4 to get x^2 -x +1/4 - 1/4 -1 =0 and now if we just factor x^2 -x +1/4 we get (x-1/2)^2 so let us sub this in to get (x-1/2)^2 = 1/4 + 1 --> (x-1/2)^2=5/4 :)

OpenStudy (anonymous):

ok thanks